an carnot engine absorbs heat at high temperature 800k and efficiency of the carnot engine is 50% what is the high temperature to increase efficiency to 80% if the low temperature kept constant ?

Grade:12th pass

2 Answers

Khimraj
5 years ago

Efficiency n = (t2-t1)/ t2 

Where t2 is the source temperature and t1 is the sink temperature  

Now when the efficiency is 50% 

0.5 = (800-t1)/800

t1 = 400K 

When the efficiency is  

0.8 = (t2-400)/t2 

The required temperature of the sink will be  

T1 = 2000k

Shubhangi Shilendra
5 years ago
Since Efficiency= 1- T2/T1,
50/100=1-T2/800
1/2=(800-T2)/800
2T2=800
T2=400K
 Now since T2 is to be kept constant and T1 is to be calculated considering 80% efficiency
so, 80/100=1-(400/T1)
4/5= (T1-400)/T1
T1=2000K